OpenStudy (anonymous):

graph the function. -2 + sin x. someone help pls.

OpenStudy (helder_edwin):

you have \(y=-2+\sin x\)

OpenStudy (helder_edwin):

do u know the graph of \(y=\sin x\)?

OpenStudy (anonymous):

the graph of sinx should intersect on pi, 2pi etc and the graph of coax should intersect the x axis at pi/2, 3pi/2 etc. Am I right?

OpenStudy (helder_edwin):

yeah, but do u know it shape?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sin x starts from 0 on the y int and cost starts from one. right?

OpenStudy (helder_edwin):

http://www.wolframalpha.com/input/?i=graph+y%3Dsin+x

OpenStudy (helder_edwin):

now the -2 outside of sin x is a vertical movement of the graph is it upward or downward?

OpenStudy (anonymous):

downward

OpenStudy (helder_edwin):

good

OpenStudy (helder_edwin):

so just move the whole graph 2 units downward do u get this?

OpenStudy (anonymous):

I moved my graph down to -2 on the y-axis and just kept the original x-int but not only going up to -1. but that seems to be wrong. =/

OpenStudy (helder_edwin):

u have to move the whole graph http://www.wolframalpha.com/input/?i=graph+y%3D-2%2Bsin+x

OpenStudy (anonymous):

yes, seems like I did it right but just didn't see it. lol. Thank u so much!

OpenStudy (helder_edwin):

u r w

OpenStudy (anonymous):

can u help me with this pls. graph: y=cos2x. I'm kind of confused with the period. I know that it is 2pie/k and I think that we have to divide this by 4. after we get the answer for that what do I have to do?

OpenStudy (helder_edwin):

ok 2 is multiplying x

OpenStudy (helder_edwin):

if u r familir with the graph of cos x then if \(x=\pi/4\) then \(2x=\pi/2\) right?

OpenStudy (helder_edwin):

sorry here it's late. http://www.wolframalpha.com/input/?i=graph+y%3Dcos+2x%2C+y%3Dcos+x

OpenStudy (anonymous):

ok thank u.